# COMI-LINGUA: Expert Annotated Large-Scale Dataset for Multitask NLP in Hindi-English Code-Mixing

### v2

@@ 4.2 Evaluation Metrics
We employ a suite of standard evaluation metrics, appropriately chosen for each task’s nature. For token-level classification tasks—LID, POS, and NER—we report Precision (P), Recall (R), and the F1-score, computed at the macro level. For the MLI task, which is a sentence-level classification problem, we adopt the same classification metrics—P, R, and F1—computed on a per-sentence basis. For MT, we use the BLEU score (Papineni et al., 2002) to evaluate the quality of translated outputs. Given the multilingual nature of our dataset, BLEU is computed separately for each output format: for English, for Romanized Hindi, and for Devanagari Hindi. This disaggregated evaluation helps assess script-specific translation quality and is especially relevant given the transliteration variability in informal code-mixed text.
@@ 4.3 Evaluation Configurations
We evaluate model performance under two distinct inference paradigms: zero-shot and one-shot (specifically, 1-shot) in-context learning. Traditional NLP tools and libraries are inherently limited to zero-shot settings, as they rely on fixed rule-based or statistical models without the capability for contextual adaptation. In contrast, LLMs are evaluated under both zero-shot and 1-shot configurations to investigate their ability to generalize from instructions alone and to leverage minimal contextual supervision, respectively.
In the zero-shot setting, the prompt includes only task-specific instructions and formatting constraints without any illustrative examples. For the 1-shot setting, we augment the prompt with a single representative example demonstrating the input-output structure of the task. This example is carefully selected to reflect typical task behavior and is kept fixed across all evaluations to maintain consistency. Detailed illustrations of both prompt configurations are provided in Appendix B.
@@ 5 Results and Observations
Tables 4 and 5 present the empirical results obtained under the two experimental configurations: zero-shot and one-shot in-context learning, respectively. It is important to note that traditional tools such as codeswitch and Microsoft LID are limited in their task coverage; consequently, results for tasks not supported by these tools are omitted from the tables.
@@
Traditional Tools vs. LLMs: The comparative analysis of traditional NLP tools and LLMs reveals clear distinctions in performance across code-mixed tasks. As shown in Table 4, traditional tools such as codeswitch and Microsoft LID demonstrate strong performance on specific tasks they were designed for, particularly POS and LID, respectively. For instance, codeswitch achieves the highest POS F1-score of 88.2, outperforming all LLMs in this task, while Microsoft LID attains a reasonable F1-score of 74.4 for LID. However, these tools exhibit significant limitations in task coverage; they do not support MLI, MT, or tasks involving complex reasoning or generation.
Open vs. Closed LLMs
@@
The performance gap between proprietary (closed) and open-source LLMs is evident across both zero-shot and few-shot settings. In zero-shot mode, closed models such as gpt-4o and claude-3.5-sonnet dominate with top-tier results in most tasks. For example, gpt-4o achieves 92.7 F1 on LID and 98.1 F1 on MLI, while claude-3.5-sonnet reaches 92.1 F1 on LID and 90.0 F1 on MLI. However, when moving to a one-shot setting, open-source models like LLaMA-3.3-instruct start closing the gap. Its performance improves significantly: LID F1 rises from 73.3 to 89.3, POS tagging reaches 84.1 (even surpassing gpt-4o), and NER climbs to 78.4. MT performance also peaks at 62.2 for English, the highest across all models.
Zero vs. One-shot Inference
@@
The transition from zero-shot to one-shot inference leads to notable performance improvements across most models and tasks. This is especially evident in complex tasks such as NER and MT, where providing one task-specific instance helps models disambiguate entities and manage code-mixed structures more effectively. For example, Claude-3.5-sonnet’s NER F1 score increases significantly from 56.7 in the zero-shot setting to 85.0 in the one-shot setting, while its for Devanagari Hindi translation improves from 31.9 to 63.4. gpt-4o similarly benefits, with NER performance rising from 60.1 to 76.0 and Devanagari improving from 31.9 to 58.26. Open models like LLaMA-3.3-instruct also see considerable gains, such as POS tagging jumping from 68.9 to 84.1 and English MT reaching 62.2. These results demonstrate that even minimal supervision through a single example can significantly enhance model performance on linguistically complex, low-resource, or code-mixed tasks. At the same time, tasks like MLI exhibit relatively modest gains, suggesting that more deterministic tasks benefit less from one-shot prompting. Overall, one-shot inference provides a practical and effective method to unlock the latent capabilities of LLMs in multilingual and code-mixed scenarios.
@@ 6 Challenges with Current LLMs
” were frequently misclassified as Hindi, reflecting a gap in script-aware language identification. Another prominent issue is sentence truncation; longer code-mixed inputs often lead to incomplete or abruptly cut-off outputs, indicating that many models struggle to preserve context over extended sequences. In addition, models such as gemini-1.5-flash and mistral-instruct displayed repetitive generation patterns, producing redundant phrases within the same response. These models also occasionally injected subjective explanations into their outputs, despite clear instructions to extract objective information—for instance, adding interpretive statements when identifying the matrix language. A further concern is that several models tended to mirror patterns from the prompt rather than perform actual analysis, indicating shallow understanding and a tendency to copy input structures. Sentences with high grammatical or script variability posed yet another barrier, where many models, especially gemini-1.5-flash and mistral-instruct, failed to generate any output at all. Overfitting to examples also emerged as a concern, particularly in one-shot settings; models like gpt-4o and command-a-03-2025 occasionally produced outputs that mimicked example structures rather than responding appropriately to the test input. This over-reliance was particularly evident in tasks such as MLI and LID, where one-shot performance slightly declined. Additionally, some models exhibited hallucination behaviors, introducing entities not present in the input, which likely stems from overgeneralization in the presence of minimal supervision (See Table 8 in the Appendix).
@@ 7 Conclusion and Future Directions
LLMs often struggle with tasks like POS tagging, NER, and MT in Hinglish code-mixed text due to limited exposure to Indian multilingual data. This results in errors such as entity mislabeling and content hallucination, especially with structurally complex and script-variable inputs. The COMI-LINGUA dataset addresses these challenges by providing high-quality, richly annotated, and task-diverse code-mixed text. Future work includes expanding the dataset and adding tasks such as normalization, sentiment analysis, Q&A, summarization, and dialogue understanding, creating a stronger foundation for training and evaluating LLMs on complex scenarios.
@@ A.3 Annotation Guidelines for All Tasks
Annotators were encouraged to flag uncertain, ambiguous, or low-quality samples for further review.
@@ Appendix B Experimental Setup
B.1 Zero-Shot Prompt
Few-Shot Prompt
@@ B.1 Zero-Shot Prompt
Twitter, YouTube, Press Releases, News (Ours) | 25,772 | R & D | Yes | 3 | |
@@
NER | Facebook (Bali et al., 2014) | 1,062 | R & D | Yes | 3 |
Twitter (Singh et al., 2018a) | 2,079 | R | Yes | 3 | |
Twitter (Singh et al., 2018c) | 3,638 | R | Yes | 2 | |
Tourism, News (Murthy et al., 2022) | 108,608 | R & D | No | 1 | |
Twitter, YouTube, Press Releases, News (Ours) | 24,913 | R & D | Yes | 3 | |
@@
Twitter, YouTube, Press Releases, News (Ours) | 24,558 | R & D | Yes | 2 | |
MLI | Twitter, Facebook (Sequiera et al., 2015) | 628 | R & D | No | 1 |
Facebook (Bali et al., 2014) | 1,062 | R & D | Yes | 3 | |
Social Media (Dhar et al., 2018) | 6,096 | R | Yes | 4 | |
Twitter, YouTube, Press Releases, News (Ours) | 25,772 | R & D | Yes | 3 |
@@
In “Yes, we belong to this place – ![]() ![]() ![]() |
Repetitive Generation | Models like gemini-1.5-flash and mistral-instruct frequently exhibit repetitive generation patterns. For instance, they may produce outputs such as: “The second tagging is more accurate as it identifies ‘this’ as a determiner and ‘last’ as a quantity. repeating similar explanations or sentence fragments within the same response. |
Subjective Additions | Instead of remaining factual, models add speculative commentary (e.g., “en: The given text is in English. The hashtag "#MadeByGoogle" is also in English. ‘E’ (English).”). |
Prompt Mimicry | gpt-4o and command-r-plus mirror example formats from the prompt, failing to adapt to new inputs and instead mimicking example structure. Based on the given text, it is written in the Hindi language. Therefore, the matrix language label for this sentence is ‘h’. |
@@
Appendix C Computation Requirement and Budget
The experiments were conducted using API-based access to state-of-the-art Large Language Models (LLMs), including gpt-4o, Command R+ (command-a-03-2025) by Cohere, and claude-3.5-sonnet. The estimated monthly costs for API usage were approximately $200 for Claude-3.5-sonnet, $150 for Cohere, and $50 for gpt-4o, resulting in a total estimated cost of $400 per month.

### v3

@@ 4.2 Evaluation Metrics
We employ a suite of standard evaluation metrics, appropriately chosen for each task’s nature. For token-level classification tasks—LID, POS, and NER—we report Precision (P), Recall (R), and the F1-score, computed at the macro level. For the MLI task, which is a sentence-level classification problem, we adopt the same classification metrics—P, R, and F1—computed on a per-sentence basis. For MT, we use the BLEU score (Papineni et al., 2002) and chrF++ score (Popović, 2015) to evaluate the quality of translated outputs. Given the multilingual nature of our dataset, BLEU and chrF++ is computed separately for each output format: , for English, , for Romanized Hindi, and , for Devanagari Hindi. This disaggregated evaluation helps assess script-specific translation quality and is especially relevant given the transliteration variability in informal code-mixed text.
@@ 4.3 Evaluation Configurations
We evaluate model performance under three distinct paradigms: zero-shot and one-shot in-context learning, and task-specific fine-tuning. Traditional NLP tools and libraries are inherently limited to zero-shot settings, as they rely on fixed rule-based or statistical models without the capability for contextual adaptation. In contrast, LLMs are evaluated under both zero-shot, one-shot and fine-tuned configurations to investigate their ability to generalize from instructions alone and to leverage minimal contextual supervision, respectively.
In the zero-shot setting, the prompt includes only task-specific instructions and formatting constraints without any illustrative examples. For the 1-shot setting, we augment the prompt with a single representative example demonstrating the input-output structure of the task. This example is carefully selected to reflect typical task behavior and is kept fixed across all evaluations to maintain consistency. For fine-tuning, we train models on task-specific training splits using formatted instruction-response pairs, allowing models to learn code-mixing patterns and task structures through parameter updates. Detailed illustrations of both prompt configurations for each task are provided in Appendix §B and for fine-tuning, detailed hyperparameters are provided in Appendix §B.1.
@@ 5 Results and Observations
Table 4 present the empirical results obtained under the two experimental configurations: zero-shot and one-shot in-context learning, respectively. It is important to note that traditional tools such as codeswitch and Microsoft LID are limited in their task coverage; results for tasks not supported by these tools are omitted from the tables.
@@
Traditional Tools vs. LLMs: The comparative analysis of traditional NLP tools and LLMs reveals clear distinctions in performance across code-mixed tasks. As shown in Table 4, traditional tools such as codeswitch and Microsoft LID demonstrate strong performance on specific tasks they were designed for, particularly POS and LID, respectively. For instance, codeswitch achieves the highest POS F1 of 88.2, outperforming all LLMs in this task, while Microsoft LID attains a reasonable F1 of 74.4 for LID. However, these tools exhibit significant limitations in task coverage; they do not support MLI, MT, or tasks involving complex reasoning or generation.
@@
Open vs. Closed LLMs The performance gap between proprietary (closed) and open-weight LLMs is evident across both zero-shot and few-shot settings. In zero-shot mode, closed models such as gpt-4o and claude-3.5-sonnet dominate with top-tier results in most tasks. For example, gpt-4o achieves 92.7 F1 on LID and 98.1 F1 on MLI, while claude-3.5-sonnet reaches 92.1 F1 on LID and 90.0 F1 on MLI. However, when moving to a one-shot setting, open-weight models like LLaMA-3.3-instruct start closing the gap. Its performance improves significantly: LID F1 rises from 73.3 to 89.3, POS tagging reaches 84.1 (even surpassing gpt-4o), and NER climbs to 78.4. MT performance also peaks at 62.2 and 74.8 for English, the highest across all models.
@@
Zero vs. One-shot Inference The transition from zero-shot to one-shot inference leads to notable performance improvements across most models and tasks. This is especially evident in complex tasks such as NER and MT, where providing one task-specific instance helps models disambiguate entities and manage code-mixed structures more effectively. For example, claude-3.5-sonnet’s NER F1 increases significantly from 56.7 in the zero-shot setting to 85.0 in the one-shot setting, while LLaMA-3.3-instruct’s improves from 55.4 to 62.2, alongside scores increasing from 71.4 to 74.8 and from 68.1 to 71.3. gpt-4o similarly benefits, with NER performance rising from 60.5 to 77.4 and improving from 32.0 to 58.4 and from 41.8 to 70.1. Open models like LLaMA-3.3-instruct also see considerable gains, such as POS tagging jumping from 68.9 to 84.1 and MT reaching 74.8. These results demonstrate that even minimal supervision through a single example can significantly enhance model performance on linguistically complex, low-resource, or code-mixed tasks. At the same time, tasks like MLI exhibit relatively modest gains, suggesting that more deterministic tasks benefit less from one-shot prompting. Overall, one-shot inference provides a practical and effective method to unlock the latent capabilities of LLMs in multilingual and code-mixed scenarios.
Fine-Tuning LLMs with COMI-LINGUA To further explore model performance beyond zero- and one-shot prompting, we fine-tuned the LLaMA-3.1-8B-Instruct, aya-expanse-8b and Qwen2.5-7B-Instruct models separately on each of the five COMI-LINGUA tasks using the respective training splits. Fine-tuning was carried out using task-specific formatted instructions, allowing the model to internalize both code-mixing patterns and structural nuances. Table 6 present the empirical results obtained under the three experimental configurations: zero-shot and one-shot in-context learning and fine-tuning respectively. The results demonstrate notable improvements across all tasks, with fine-tuned models outperforming traditional tools, open-weight baselines, and in some cases closed-weight LLMs.
Fine-tuning vs. Prompting Approaches The fine-tuning results from Tables 6 and 7 demonstrate substantial performance gains over both zero-shot and one-shot inferencing approaches. Fine-tuned models consistently outperform traditional tools and achieve competitive or superior results compared to closed-weight LLMs across all tasks. Particularly notable improvements are observed in NER (30-40% gains over prompting approaches) and consistent high performance in MLI (>95% F1 across all fine-tuned models). In the MT task, the model yielded BLEU scores of 56.1 for , 66.6 for and 73.5 for . Correspondingly, the chrF++ scores were 78.7 for , 85.9 for and 86.2 for MT.
These results highlight the strength of supervised fine-tuning on high-quality, diverse code-mixed data, as provided by COMI-LINGUA. Unlike zero- or one-shot setups, which depend heavily on prompt engineering and model prior knowledge, fine-tuning allows the model to generalize deeper linguistic patterns and task-specific strategies.
@@ 6 Challenges with Current LLMs
” were frequently misclassified as Hindi, reflecting a gap in script-aware language identification. Another prominent issue is sentence truncation; longer code-mixed inputs often lead to incomplete or abruptly cut-off outputs, indicating that many models struggle to preserve context over extended sequences. Models such as gemini-1.5-flash and mistral-instruct displayed repetitive generation patterns, producing redundant phrases within the same response. These models also occasionally injected subjective explanations into their outputs, despite clear instructions to extract objective information—for instance, adding interpretive statements when identifying the matrix language. Several models tended to mirror patterns from the prompt rather than perform actual analysis, indicating shallow understanding. Sentences with high grammatical or script variability posed yet another barrier, where many models, especially gemini-1.5-flash and mistral-instruct, failed to generate any output at all. Overfitting to examples also emerged as a concern, particularly in one-shot settings; models like gpt-4o and command-a-03-2025 occasionally produced outputs that mimicked example structures rather than responding appropriately to the test input. This over-reliance was particularly evident in tasks such as MLI and LID, where one-shot performance slightly declined. Additionally, models hallucinated non-existent entities, suggesting overgeneralization from minimal supervision. (See Table 10 in the Appendix).
Beyond these general limitations, our analysis of smaller models (7–8B parameters) uncovered some failure patterns; for MT task, Qwen2.5-7B-Instruct inappropriately provided empty outputs with structured labels such as “Unit 1: English, Unit 2: Romanized Hindi, Unit 3: Devanagari Hindi” instead of providing actual translations. In NER tasks, aya-expanse-8b misclassified punctuation marks, tagging “(” as Opening parenthesis and “)” as Closing parenthesis rather than using standard entity categories and labelling as ‘X’. POS tagging revealed more hallucination patterns, with LLaMA-3.1-8B-Instruct generating repetitive sequences like “VERB NOUN PROPN NOUN NOUN NOUN NOUN” for multiple instances. More concerning was the tendency of these models to output code snippets instead of task responses —LLaMA-3.1-8B-Instruct and Qwen2.5-7B-Instruct output Python import statements and function templates, rather than returning direct predictions as:
import nltk
from nltk import pos_tag
# Download the required NLTK data
nltk.download(‘perceptron_tagger’)
nltk.download(‘punkt’)
Similar code generation patterns emerged across tasks, with the model providing import re and langdetect modules for LID task rather than providing the actual labels. Entity hallucination was prevalent in Qwen2.5-7B-Instruct, which generated anomalous labels like (Live India) as ‘X X X X X X’ and inappropriately tagged terms such as “” as HASHTAG entity. MT task suffered from incomplete generation, with outputs abruptly ending mid-sentence, as observed in Romanized Hindi Translation: “Madras Hāikōrt ne dāk vibhāg ko.”(See Table 11 in the Appendix). These systematic failures across all three smaller models highlight the importance of robust fine-tuning and careful prompt engineering when deploying compact LLMs for complex multilingual tasks.
@@ 7 Conclusion and Future Directions
LLMs often struggle with tasks like POS tagging, NER, and MT in code-mixed Hindi-English due to their lack of exposure to Indian multilingual data. Errors such as mislabeling entities or hallucinating content arise from limited training on structurally complex and script-variable inputs. The COMI-LINGUA dataset addresses these issues by providing high-quality, task-diverse, and richly annotated code-mixed text. Fine-tuning on this dataset enables models to better handle linguistic ambiguity, reduce overfitting, and improve reliability across tasks. Its inclusion of contextual examples and diverse sources—like social media and news—enhances the models’ ability to generalize across formal and informal registers, while iterative refinement through active learning ensures sustained performance gains.
@@
4.
Computational Accessibility: While fine-tuning shows substantial improvements, computational requirements and the need for substantial training data may limit accessibility for resource-constrained settings.
@@
Acknowledgments
This work is supported by the Anusandhan National Research Foundation (ANRF), India, through the project titled “Curating and Constructing Benchmarks and Development of ML Models for Low-Level NLP Tasks in Hindi-English Code-Mixing”. The authors express their gratitude to Diksha, Ronakpuri Goswami, Mahesh Kumar, Rahul Gadhvi, Yash Chopra, Mahavir Patil, Vaidahi Patel and Ashish Singh for their invaluable support with dataset annotations. We also extend our thanks to Sailesh Panda, Isha Narang and Prathamesh Shanbhag for their assistance in reveiwing the manuscript and providing feedback. Himanshu Beniwal is supported by the Prime Minister Research Fellowship (PMRF ID-1702154), India.
@@ A.3 Annotation Guidelines for All Tasks
Annotators were encouraged to flag uncertain, and ambiguous samples for further review.
@@
Quality Control & Training
Annotators periodically used gold-standard examples to ensure continued alignment throughout the annotation process.
Periodic sample checks provided feedback and helped uphold annotation standards.
An independent reviewer regularly flagged low-quality annotations for re-annotation by the original annotators.
Conflict Resolution Strategy
Consolidated annotation criteria: For model training and evaluation, only annotations with agreement from at least two out of three annotators were retained, ensuring reliability.
Iterative refinement: Disagreement patterns were analyzed to identify common sources of confusion, leading to guideline refinements and additional training for annotators.
@@ Appendix B Experimental Setup
Zero-shot LID Prompt
One-shot LID Prompt
Zero-shot MLI Prompt
One-shot MLI Prompt
Zero-shot NER Prompt
One-shot NER Prompt
Zero-shot POS Tagging Prompt
One-shot POS Tagging Prompt
Zero-shot MT Prompt
One-shot MT Prompt
B.1 Fine-Tuning Hyperparameters
The optimization process focused on fine-tuning four key hyperparameters, with the goal of balancing the refinement of essential parameters while minimizing unnecessary adjustments to those already well-suited for the task. Batch size, number of epochs, weight decay, and learning rate were selected due to their direct and substantial impact on model performance, stability, and generalization.
Core Training Parameters:
Epochs: 3
Batch size: 4 per device with gradient accumulation steps of 8 (effective batch size: 32)
Learning rate: 2e-4 with cosine scheduler and warmup ratio of 0.1
Weight decay: 0.01
LoRA Configuration: LoRA with rank 32, alpha 64, dropout 0.1.
Instruction Format: All tasks used task-specific instruction templates with examples, following the format:
Instruction: [task description]
Sentence: [input]
Output: [expected output]
Appendix C Computation Requirement and Budget
The experiments were conducted using API-based access to state-of-the-art Large Language Models (LLMs), including gpt-4o, Command R+ (command-a-03-2025) by Cohere, and claude-3.5-sonnet. The estimated monthly costs for API usage were approximately $200 for claude-3.5-sonnet, $150 for Cohere, and $50 for gpt-4o, resulting in a total estimated cost of $400 per month. For computational infrastructure, experiments were carried out on four NVIDIA Tesla V100 32 GB GPUs, with an estimated cost of $7,192.00 per month based on Google Cloud Platform (GCP) 202020https://cloud.google.com/products/calculator Calculator pricing.
@@ Appendix C Computation Requirement and Budget
Twitter, YouTube, Press Releases, News (Ours) | 25,773 | R & D | Yes | 3 | |
MLI | Twitter, Facebook (Sequiera et al., 2015) | 628 | R & D | No | 1 |
Facebook (Bali et al., 2014) | 1,062 | R & D | Yes | 3 | |
Social Media (Dhar et al., 2018) | 6,096 | R | Yes | 4 | |
Twitter, YouTube, Press Releases, News (Ours) | 25,773 | R & D | Yes | 3 | |
NER | Facebook (Bali et al., 2014) | 1,062 | R & D | Yes | 3 |
Twitter (Singh et al., 2018a) | 2,079 | R | Yes | 3 | |
Twitter (Singh et al., 2018c) | 3,638 | R | Yes | 2 | |
Tourism, News (Murthy et al., 2022) | 108,608 | R & D | No | 1 | |
Twitter, YouTube, Press Releases, News (Ours) | 24,913 | R & D | Yes | 3 | |
@@
Twitter, YouTube, Press Releases, News (Ours) | 24,558 | R & D | Yes | 2 |
@@
In ‘Yes, we belong to this place – ![]() ![]() ![]() |
Repetitive Generation | Models like gemini-1.5-flash and mistral-instruct frequently exhibit repetitive generation patterns. For instance, they may produce outputs such as: ‘The second tagging is more accurate as it identifies ‘this’ as a determiner and ‘last’ as a quantity’. repeating similar explanations or sentence fragments within the same response. |
Subjective Additions | Instead of remaining factual, models add speculative commentary (e.g., ‘en: The given text is in English. The hashtag ‘#MadeByGoogle’ is also in English. ‘E’ (English).’). |
Prompt Mimicry | gpt-4o and command-r-plus mirror example formats from the prompt, failing to adapt to new inputs and instead mimicking example structure. ‘Based on the given text, it is written in the Hindi language. Therefore, the matrix language label for this sentence is h’. |
@@
Response Flaw Type | Example Behavior or Observation |
Output Blank or Missing | The Qwen2.5-7B-Instruct and LLaMA-3.1-8B-Instruct models frequently returned blank or no output for complex sentences across all tasks, particularly in zero-shot settings when encountering noisy or script-mixed inputs. |
Task Instruction Violation |
Models like aya-expanse-8b and LLaMA-3.1-8B-Instruct generated Python code snippets in some task outputs:
‘# Split the input text into words based on spaces words = text.split() # Initialize an empty list for words tagged_words = [] # Iterate over each word in input text given for word in words: # Check if the word is in the Hindi’ for LID tasks, failing to follow prompt formatting requirements and giving incorrect output. |
Subjective Commentary Injection | All Models added unnecessary disclaimers like ‘Please note that the output is a best-effort attempt and might not be 100% accurate due to the complexity of Hinglish language’ and ‘Please let me know if you need any further assistance!’ instead of providing direct outputs. |
Language Script Confusion |
aya-expanse-8b exhibited severe input-output disconnection and incorrect script identification in MT and MLI tasks. For MT, an input about Bangladesh football team yielded an unrelated output discussing Republic TV and Arnab Goswami in “Urdu” script. For MLI, the input ‘![]() |
Entity Type Misclassification | All Models frequently misclassified entities in code-mixed contexts, such as tagging ‘Uttar Pradesh’ as separate mismatched entities (‘Uttar’:ORGANISATION, ‘Pradesh’:PLACE) instead of the correct unified LOCATION labels (‘Uttar’:LOCATION, ‘Pradesh’:LOCATION), demonstrating poor understanding of entity boundaries. |
Hallucination |
aya-expanse-8b in one-shot POS tagging created its own input sentence
Inconsistent Output Formats | LLaMA-3.1-8B-Instruct and Qwen2.5-7B-Instruct provided inconsistent MLI labels like ‘Mixed’, ‘Code-mixed language: Hindi-English’ instead of ‘hi’ or ‘en’, for some inputs, showing format instability across evaluation instances. |
Multi-language Script Errors | For zero-shot MT, Qwen2.5-7B-Instruct generated Romanized Hindi in Arabic-Urdu, Bangla-English script instead of Latin script, completely misunderstanding and hallucinating in some input instances. |